OpenStudy (anonymous):

What is the volume of a rectangular based pyramid with a height of eighteen meters and base dimensions of 5 meters and ten meters? one hundred cubic meters three hundred cubic meters nine hundred cubic meters two thousand seven hundred cubic meters

jhonyy9 (jhonyy9):

so do you know formula for volum ? is area of base time height divide 3 hope this will help you

jhonyy9 (jhonyy9):

what you wann with pi here ?

jhonyy9 (jhonyy9):

so area of base will be 5*10=50 volum = 50*18/3 = ?

jhonyy9 (jhonyy9):

so V=50*6=300 m^3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Yes, that's what I got didn't know if i was correct or not, thank you. :)

jhonyy9 (jhonyy9):

was my pleasure bye